An asset used in a 4-year project falls in the 5-year MACRS class (MACRS Table) for tax purposes. The asset has an acquisition cost of $11,340,000 and will be sold for $2,520,000 at the end of the project.

  

If the tax rate is 24 percent, what is the aftertax salvage value of the asset?
Answer #1

book value of asset = cost - depreciation

= 11340000 - 11340000 * (0.2 + 0.32 + 0.1920 + 0.1152)

= 1959552

tax on sale = (sale price - book value) * (tax rate)

= (2520000 - 1959552) * (0.24)

= 134507.52

after tax salvage value

= price - tax on sale

= 2520000 - 134507.52

= 2385492.48